# Inferring Latent Temporal Sparse Coordination Graph for Multi-Agent Reinforcement Learning

### v1

@@ V-B1 Trajectory Graph Learning vs One-Step Observations
Figure 7: Evaluate the effect of the different weights of graph loss.
V-B2 Latent Temporal Sparse Graph Learning strategies
@@
  • LTS-CG with both Predict-Future and Infer-Present characteristics, where we consider both aspects simultaneously.
The final performance is assessed on the 8m_vs_9m, and 3s5z maps and the results are presented in Fig. 6. The ablation study revealed several important findings. Firstly, regardless of whether we include the Predict-Future, the Infer-Present, or both characteristics, the performance was consistently better than not having anyone. This highlights the importance of having these characteristics in enhancing the learning of the inter-agent graph and improving cooperative behaviour. Moreover, on the 8m_vs_9m map, with the Predict-Future characteristic outperformed the other settings. One possible reason for this observation is that the agents in this map are homogeneous, sharing similar characteristics. Knowing the next time observation benefits the overall cooperation among the agents. In contrast, in the 3s5z map agents are heterogeneous. Utilizing observations from different agents to infer the current state proves beneficial for learning in this scenario. Although using both characteristics simultaneously introduces more parameters and slightly slower convergence, promising results are obtained after approximately 2 million time steps, showcasing the effectiveness of leveraging both characteristics for improved cooperative multi-agent learning. This ablation study confirms the significance of two characteristics in LTS-CG for learning meaningful graphs to help agents cooperate.
V-B3 Weight of graph loss
@@
V-C Discussion
In this discussion, we underscore the contrasts between our proposed approach and CG-based methods (e.g., DCG [10], SOP-CG [13], and CASEC [14]), as well as with the method where the graph serves for information exchange (e.g., DICG [12], G2ANet [21]).
[…]
DCG $11.63\pm 0.64$ $O(A^{2}N^{2})$
SOP-CG $19.46\pm 0.80$ $O(A^{2}N^{2})$
CASEC $10.12\pm 0.51$ $O(A^{2}N^{2})$
LTS-CG $8.84\pm 0.49$ $O(TN^{2})$
TABLE IV: Time consumption on map 10m_vs_11m and complexity analysis with CG-based methods.
V-C1 Coordination graphs (CG) methods
The CG is denoted as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ where $\mathcal{V}$ is agent/node set and $\mathcal{E}$ is the edge set between agents [8, 41]. These CG-based methods factorize the Q-function into utility functions $q^{i}$ and payoff functions $q^{ij}$ as follows:
$Q_{tot}\!\left(s_{t},\boldsymbol{a}\right)\!=\!\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}q^{i}\left(a^{i}\!\mid\!s_{t}\right)\!+\!\frac{1}{|\mathcal{E}|}\!\sum_{\{i,j\}\in\mathcal{E}}\!\!\!q^{ij}\!\left(a^{i},a^{j}\!\mid\!s_{t}\right).$ (18)
However, due to the large number of action pairs represented by the payoff functions $q^{ij}$, these methods face challenges of high computational complexity. To alleviate this concern, various approaches have been employed, such as low-rank approximation in DCG [10], construction of polynomial-time CG in SOP-CG [13], and dropping edges using variance payoff functions in CASEC [14]. Despite these efforts, the inherent high complexity of CG-based methods, as high as $O(A^{2}N^{2})$, limits their applicability in large-scale scenarios.
By contrast, the complexities of LTS-CG is $O(TN^{2})$, where $N$ is the number of agents, and $T$ is the length of trajectories. For instance, consider experiments conducted on the map 10m_vs_11m, involving 10 agents and 17 actions, resulting in 289 potential action-pair within $q^{ij}$. Table IV illustrates the time requirements of various methods, revealing that CG-based approaches demand more time to complete $1m$ time steps than our method. Particularly noteworthy is the significantly higher time consumption of SOP-CG and the remarkably elevated GPU usage of CASEC (25.85 GB), contrasted with DCG (2.29 GB), SOP-CG (4.00 GB), and our approach (4.13 GB). Moreover, it’s important to highlight that all CG-based methods failed to conclude the experiment on the 27_vs_30m map, involving the calculation of $129636^{2}=$ 1296 action-pair for every two agents (the complexity now is $O(1296\times N^{2})$). In contrast, for our proposed LTS-CG method, even though we set the length of trajectory for graph learning to the episode limit, the computational complexity remains a modest $O(180\times N^{2})$. In the actual experiment, this trajectory length typically falls below the episode limit. Thus, LTS-CG delivers competitive performance while maintaining a reasonable level of computational efficiency.
Figure 8: Visualization of latent agent-pair matrices learned by DICG and LTS-CG on map 8m_vs_9m.
V-C2 Graphs for information exchange
Current methods using graphs for information exchange rely on attention mechanisms to calculate edge weight. Like DICG [12], these methods learn a weighted fully-connected graph, which inevitably transfers redundant or even detrimental information between agents, as illustrated in Fig.8. This drawback impedes individual agents’ capacity to acquire effective policies. In contrast, our method generates a sparser graph focusing on the most relevant relationships among agents, leading to more efficient and effective cooperation.
Another existing method is G2ANet [21], which uses both hard and soft attention to generate a sparse graph for MARL. However, our method offers two distinct advantages over G2ANet: (1) G2ANet focuses solely on ongoing information, while we leverage the complete observation trajectories to capture the latent relationships between agents, shown in Sec.V-B1. (2) Unlike G2ANet’s arbitrary edge reduction using the attention mechanism, our approach harnesses agents’ experience and current information to acquire a meaningful graph. This graph exhibits two critical attributes, the efficacy of which has been demonstrated through Sec.V-B2.

### v2

@@ V-C1 Trajectory Graph Learning vs One-Step Observations
V-C2 The Necessity of Sampling from Attention Matrix
We further investigated whether sampling the graph from the attention matrix (sparse graph) is more effective than not sampling (dense graph). In the latter case, the attention matrix is directly used as the adjacency matrix, resulting in a fully connected graph. These studies were performed on the 3s5z and 10m_vs_11m maps. The results are presented in Fig. 8.
Our method outperforms the fully connected graph approach, where the attention matrix is used directly. This suggests that relying attention-based matrix is insufficient for optimal performance. One potential reason for the lower performance is the excessive exchange of messages between agents. While communication aims to enhance coordination, the large volume of irrelevant messages can overwhelm agents and distract them from making optimal decisions.
In contrast, our method’s sparse graph mitigates this issue by restricting message passing to the most relevant agent pairs, reducing unnecessary information flow. This allows agents to focus on the most critical knowledge for decision-making. Furthermore, by treating the attention matrix as a distribution and sampling the graph from it, LTS-CG captures the inherent uncertainty in dynamic environments more effectively, leading to richer representations of agent cooperation and better adaptability over time.
V-C3 Latent Temporal Sparse Graph Learning strategies
@@
  • LTS-CG with both Predict-Future and Infer-Present characteristics.
The final performance is assessed on the 8m_vs_9m, 3s5z maps of SMAC and the Gather scenario. The results are presented in Fig. 9. The ablation study revealed several important findings. Firstly, regardless of whether we include the Predict-Future, the Infer-Present, or both characteristics, the performance was consistently better than not having anyone. This trend is consistent across different environments (SMAC and Gather), highlighting that each characteristic independently enhances agent coordination and overall performance. Moreover, since SMAC is known for its complexity and dynamic nature, our method’s performance, incorporating both $\mathcal{L}_{pre}$ and $\mathcal{L}_{inf}$, exhibits some variation across different maps. This dynamic performance underscores that different maps may require different levels of emphasis on prediction and inference. To gain deeper insights into these dynamics, we extended our investigation to the Gather environment, which offers a different set of challenges and complexities. The results in this environment confirm the generalizability of our findings: $\mathcal{L}_{pre}$ is not redundant but plays a crucial role in improving performance when used in conjunction with $\mathcal{L}_{inf}$ The synergy between these losses ensures that our method can effectively capture temporal dependencies and uncertainties in agent relationships, leading to superior outcomes in multi-agent coordination tasks. This ablation study confirms the significance of two characteristics in LTS-CG for learning meaningful graphs to help agents cooperate.
Figure 10: Evaluate the effect of the different weights of $\mathcal{L}_{g}$.
V-C4 Weight of Graph Loss
@@
Since $\mathcal{L}_{g}$ comprises two components, $\mathcal{L}_{pre}$ (predict-future) and $\mathcal{L}_{inf}$ (infer-present), we further investigated their individual contributions by testing different weight combinations $\{0,1,5,10,50\}$ for each, with 5 independent runs for each setting. The results, shown in Fig. 11, depict the average test win rate under various weight configurations.
From this analysis, we observed the following key points: (1) Even a minimal inclusion of either $\mathcal{L}_{pre}$ or $\mathcal{L}_{inf}$ (i.e., weights greater than 0) consistently improves the test win rates compared to their exclusion. This finding validates the effectiveness of introducing the Predict-Future and Infer-Present mechanisms, leading to more meaningful graph construction and improved agent coordination. (2) As the weights of $\mathcal{L}_{pre}$ and $\mathcal{L}_{inf}$ increase from 0 to 10, the performance improves steadily. However, as the weights further increase to 50, the performance starts to degrade. This trend suggests that a moderate weighting of these losses is beneficial, while overly large weights may negatively affect performance by overemphasizing certain aspects of the learning process. (3) The best performance is achieved when both the weights of $\mathcal{L}_{pre}$ and $\mathcal{L}_{inf}$ are set to 10, as this configuration yields the highest test win rate, indicating an optimal balance between the two loss terms.
@@
Figure 11: Heatmap illustrating the average test win rate on the 8m_vs_9m map in the SMAC environment under different configurations of the $\mathcal{L}_{pre}$ and $\mathcal{L}_{inf}$ weights.
Figure 12: A case study on the StarCraft II benchmark map 3s5z, featuring 3 Stalkers and 5 Zealots. The top row displays screenshots from the actual gameplay replay, while the bottom row illustrates the corresponding attention matrices and final sparse matrices.
V-D Case Study
In this case study, we visualize the attention and sparse matrices alongside the actual game replay to demonstrate the interpretability of our model, shown in Fig.12. It highlights the most critical interactions among agents at different stages of the game:
  • Scenario (a): At the beginning of the game, all agents exhibit high attention values towards each other (notably in the last column and row of the attention matrix), underscoring the importance of initial coordination. Even at this early stage, the sparse matrix begins reducing edges, refining communication to focus on key interactions.
  • Scenario (b): During the combat phase, where three Zealots are actively engaged while other agents remain on standby, the attention matrix reveals two distinct blocks that correspond to the two separate groups of agents. The sparse matrix emphasizes the importance of within-group communication over interactions between the groups at this point in the game.
  • Scenario (c): After the elimination of agents 4 and 5, the attention matrix shows reduced intensity in the rows and columns corresponding to these agents. The sparse matrix further prunes edges related to the eliminated agents, effectively modeling the decreased necessity for their participation in the communication network.
These visualizations of the attention matrix help illustrate how LTS-CG dynamically captures the most relevant relationships among agents, contributing to a clearer understanding of the method.
Method Graph type
Sample
Edge
Data used
Graph Calculation
Time Complexity
QMIX $\times$ $\times$ $\times$ $\times$
DCG Complete $\times$ One-step $O(A^{2}N^{2})$
DICG Complete $\times$ One-step $O(KN^{2})$
SOP-CG Sparse $\times$ One-step $O(A^{2}N^{2})$
CASEC Sparse $\times$ One-step $O(A^{2}N^{2})$
LTS-CG Sparse Yes Trajectories $O(TN^{2})$
TABLE I: Comparison of different experiment methods in terms of graph type, edge sampling, data used for learning the graph, and graph calculation time complexity.
Method 1k steps time (s) 1m steps time (h) GPU Memory
QMIX $15.21\pm 2.48$ $2.7\pm 0.41$ 1.32 GB
DCG $32.50\pm 1.71$ $11.25\pm 1.47$ 1.59 GB
DICG $20.12\pm 2.76$ $7.32\pm 1.69$ 1.60 GB
CASEC $28.50\pm 4.65$ $9.46\pm 1.82$ 6.38 GB
SOP-CG $24.22\pm 5.68$ $13.90\pm 0.56$ 2.45 GB
LTS-CG $20.76\pm 2.47$ $5.64\pm 1.53$ 3.17 GB
TABLE II: Running time and GPU consumption on 8m.
Method 1k steps time (s) 1m steps time (h) GPU Memory
QMIX $20.13\pm 3.59$ $6.79\pm 0.37$ 1.50 GB
DCG $33.57\pm 4.65$ $11.63\pm 0.64$ 2.37 GB
DICG $20.74\pm 4.37$ $7.81\pm 0.46$ 2.03 GB
CASEC $30.50\pm 2.03$ $10.12\pm 0.51$ 10.64 GB
SOP-CG $35.46\pm 3.62$ $19.46\pm 0.80$ 4.21 GB
LTS-CG $22.68\pm 3.89$ $8.84\pm 0.49$ 4.45 GB
TABLE III: Running time and GPU consumption on 10m_vs_11m.
Method 1k steps time (s) 1m steps time (h) GPU Memory
QMIX $26.28\pm 4.58$ $7.30\pm 0.72$ 1.93 GB
DCG $43.67\pm 5.73$ $18.83\pm 0.58$ 13.35 GB
DICG $27.79\pm 6.65$ $8.94\pm 0.68$ 3.39 GB
CASEC / / Out of 48GB GPU
SOP-CG $516.04\pm 10.76$ More than 7 days 25.42 GB
LTS-CG $31.73\pm 2.89$ $10.37\pm 0.53$ 11.91 GB
TABLE IV: Running time and GPU consumption on 25m.
V-E Discussion
This section first summarizes the compared methods in terms of graph type, edge sampling, data used for learning the graph, and graph calculation time complexity, as shown in Tab.I. Next, we highlight the key theoretical differences between our LTS-CG method and existing graph-based MARL approaches.
V-E1 Graph as Coordination Graph vs. Graph for Message Passing
Existing methods like DCG [10], SOP-CG [13], and CASEC [14] explicitly model coordination between agents using a coordination graph (CG). The graph represents action pair coordination, where the Q-function is factorized into utility functions $q^{i}$ and payoff functions $q^{ij}$ as Eq. (1). This explicit coordination allows for direct evaluation of the coordination quality but incurs a high computational cost, $O(A^{2}N^{2})$, due to the large number of action pairs required for the payoff functions $q^{ij}$. In contrast, methods like DICG [12], and LTS-CG infer implicit graphs that facilitate knowledge sharing during policy learning, bypassing direct action pair calculations. This approach significantly reduces computational complexity. For example, LTS-CG operates with a complexity of $O(TN^{2})$, where $T$ is the trajectory length.
We present the detailed running time and GPU consumption for the compared methods on the 8m (Tab. II), 10m_vs_11m (Tab. III), and 25m (Tab. IV) maps from SMAC. As shown in these tables, when the number of agents increases from 8 to 25, our LTS-CG method maintains acceptable computational resource consumption. In contrast, CASEC exceeds the 48GB GPU memory limit, and SOP-CG could not complete 1 million steps within one week on the 25m map. These findings demonstrate that LTS-CG scales efficiently with larger agent counts, offering a more practical solution for complex multi-agent scenarios compared to other graph-based methods, especially in terms of computational resources and runtime efficiency.
V-E2 Sampling Graphs from Attention Matrix vs. Direct Use of Attention Matrix
Unlike the method DICG[12], which directly use the attention matrix as the graph, LTS-CG introduces a novel sampling approach. By treating the attention matrix as a distribution and sampling graphs from it, LTS-CG effectively captures the uncertainty inherent in dynamic environments. This sampling process allows for richer and more adaptable representations of agent cooperation, as the graphs evolve over time based on the sampled attention weights. Consequently, this approach enhances agent adaptability to changing environments, resulting in improved performance and coordination flexibility.
V-E3 Trajectories vs. One-step Observation
A key distinction in LTS-CG is the use of observation trajectories to generate the agent-pair probability matrix, rather than relying on single-step observations. We posit that observation trajectories provide a more comprehensive view of the temporal dynamics of agent interactions, leading to more accurate and meaningful graph representations. Empirical results in Sec. V-C1 support the validity of this assumption, demonstrating the effectiveness of trajectory-based graph construction.
V-E4 Further Learning the Graph Characteristics
Many existing graph-based methods (e.g., DCG[10], DICG[12]) rely primarily on attention mechanisms to infer coordination graphs but often lack additional regularization techniques, which can lead to arbitrary or less informative edges. LTS-CG addresses this issue by introducing two distinctive components: Predict-Future and Infer-Present. These mechanisms allow agents to anticipate future states and optimize their current coordination with limited data, respectively. By incorporating these features, LTS-CG constructs graphs that are not only spatially but also temporally optimized, leading to more meaningful and informed cooperation between agents. This enhanced graph learning, combined with regularization, results in significantly better collaboration and performance across complex, multi-agent environments.